OpenStudy (anonymous):

3X-4Y=6 6X-8Y=10 the system shown had ____ solution(s).

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (anonymous):

i think i got it? what did u get @daisylovesyou?

Directrix (directrix):

1) 3X-4Y=6 2) 6X-8Y=10 ---------------- Suggestion: Multiple every member of equation 1) by (-2) and add the "new" equation 1 to equation 2) 1) 3X-4Y=6 --> by (-2) --> -6X + 8Y = -2 2) 6X-8Y=10 1') -6X + 8Y = -2 2) 6X -8Y= 10 ------------------ Your task: Add the two equations. @daisylovesyou

OpenStudy (anonymous):

You also can divide the 2nd equation by 2